# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

### v1

@@ 4.2 Analysis of Tracking the Optimal Solution Trajectory of the Maximization Component
In the first step of examining the trajectory of the optimal solution in the maximization component, we need to discuss two important properties outlined in Lemmas 4.3 and 4.4, which have already been proven in [4] within the context of the SP problem. Here, we present the lemmas and their proofs for completeness.
@@
The solution map $y^{*}_{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$ is Lipschitz continuous. Suppose Assumption 2.1, 2.2 hold and $\theta^{*}(x)$ is Lipschitz continuous map with constant $\mathbf{L}_{\theta}$. In particular, for any $x,\overline{x}\in\mathcal{X}$
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq\frac{L_{y\mu}}{\mu}\left\|x-\overline{x}\right\|$ (21)
@@
$(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{T}\big{[}\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\big{]}+\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq 0.$ (22)
@@
$(y-y^{*}_{\mu}(\overline{x}))^{T}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))\leq 0,$ (23)
$(y-y^{*}_{\mu}(x))^{T}\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\leq 0.$ (24)
@@
$(y^{*}_{\mu}(x)-y^{*}_{\mu}(\overline{x}))^{T}\Big{(}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\Big{)}\leq 0.$ (25)
@@
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{T}\Big{(}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(\overline{x}))\Big{)}$ (26)
@@
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{T}\Big{(}\nabla_{y}\Phi_{\mu}(\overline{x},\theta^{*}(\overline{x}),y^{*}_{\mu}(\overline{x}))-\nabla_{y}\Phi_{\mu}(x,\theta^{*}(x),y^{*}_{\mu}(\overline{x}))\Big{)}$ (27)
@@
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|(L_{yx}^{\Phi}\left\|x-\overline{x}\right\|+L_{y\theta}^{\Phi}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|)$ (28)
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq\frac{L_{yx}^{\Phi}}{\mu}\left\|x-\overline{x}\right\|+\frac{L_{y\theta}^{\Phi}}{\mu}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|$ (29)
@@
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq(\frac{L_{yx}^{\Phi}+\mathbf{L}_{\theta}L_{y\theta}^{\Phi}}{\mu})\left\|x-\overline{x}\right\|$
@@
Suppose Assumption 2.1, 2.2 hold. The function $f_{\mu}(\cdot)$ is differentiable on an open set containing $\mathcal{X}$ and $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$ where $y^{*}_{\mu}(x)\triangleq\operatorname*{\arg\!\min}_{y\in\mathcal{Y}}\mathcal{L}_{\mu}(x,y)$. Moreover, $f_{\mu}$ has a Lipschitz continuous gradient with constant $\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu}$.
@@
From Danskins’s theorem [3] one can obtain $f_{\mu}(\cdot)$ is differentiable and $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$. Therefore, we have
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|=\left\|\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))-\nabla_{x}\mathcal{L}(x^{\prime},y^{*}_{\mu}(x^{\prime}))\right\|$ (30)
@@
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-x^{\prime}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\left\|y^{*}_{\mu}(x)-y^{*}_{\mu}(x^{\prime})\right\|$ (31)
@@
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-x^{\prime}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\frac{L_{y\mu}}{\mu}\left\|x-x^{\prime}\right\|$ (32)
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\left\|x-x^{\prime}\right\|$ (33)
Next, we aim to evaluate the accuracy of the iterative solution $y_{k}$ in approximating the optimal solution $y^{*}_{\mu}(x_{k})$ by bounding the error between them.
@@
Suppose Assumptions 2.1, 2.2 hold. Considering that $\{(x_{k},y_{k})\}_{k\geq 0}$ be the sequence generated by the Algorithm. Let $\sigma_{k}=\sigma\leq\frac{2}{L_{yy}^{\Phi}+2\mu}$. Then for any $k\geq 0$,
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)},$ (34)
@@
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|\leq\frac{L^{\Phi}_{y\theta}}{\mu}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)}.$ (35)
@@
$\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma\|G_{k-1}^{y}-\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k-1})\|,$ (36)
@@
$\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}\|\theta_{k}-\theta^{*}(x_{k})\|.$ (37)
@@
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}(\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|x_{i}-x_{i+1}\right\|).$ (38)
@@
where $\Gamma_{1}\triangleq\frac{1}{1-\rho_{D}}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|D_{\mathcal{X}}+\frac{\sigma L_{y\theta}^{\Phi}\beta D_{\mathcal{X}}}{(1-\rho_{D})(1-\beta)}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$ and $\Gamma_{2}\triangleq\frac{\gamma L_{y\mu}}{\mu(1-\rho_{D})}+\frac{\sigma\gamma\beta L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{(1-\beta)(1-\rho_{D})}=\mathcal{O}\left(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma\right)$.
@@
$\quad+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)},$ (41)
where in the last inequality we applied Lemma 4.3.
@@
$\quad+\rho^{k-i}_{D}\gamma\mathbf{L}_{\theta}\sum_{j=0}^{i-1}\beta^{i-j}\left\|s_{j}-x_{j}\right\|\Big{)},$ (42)
@@
Moreover, using Lemma 8.1 one can verify that for any $K\geq 1$,
@@
Finally, note that $\mathbf{L}_{\theta}=\mathcal{O}(\kappa_{g})$, $L_{y\mu}=\mathcal{O}(\kappa_{g})$, $1-\rho_{D}=\mathcal{O}(\mu)$, and $1-\beta=\mathcal{O}(1/\kappa_{g})$.
@@
Suppose Assumption 2.1, 2.2 hold. For any $k\geq 0$ we have
@@
where $\Gamma_{3}\triangleq\frac{2\Gamma_{2}}{1-\rho}+\frac{L_{y\mu}\gamma}{(1-\rho)\mu}$, and $\Gamma_{1},\Gamma_{2}$ are defined in Lemma 4.6 and $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$.
@@
Using the triangle inequality one can observe that for any $k\geq 0$,
$\left\|y_{k+1}-y_{k}\right\|\leq\left\|y_{k+1}-y^{*}_{\mu}(x_{k+1})\right\|+\left\|y^{*}_{\mu}(x_{k})-y_{k}\right\|+\left\|y^{*}_{\mu}(x_{k+1})-y^{*}_{\mu}(x_{k})\right\|$ (45)
Using Lemma 4.3 and 4.6 we can obtain
@@
Moreover, the second part of the lemma can be proved with a similar discussion as in the proof of Lemma 4.6 by employing the results in Lemma 8.1.

### v3

@@ 4.2 Analysis of Tracking the Optimal Solution Trajectory of the Maximization Component
In this section, we examine the properties of the regularized optimal solution in the maximization component of (5) as discussed in Section 3. In particular, we show that the regularized optimal solution $y^{*}_{\mu}$ is Lipschitz continuous, and the function $f_{\mu}$ has a Lipschitz continuous gradient, in Lemmas 4.3 and 4.4, respectively. Although similar results have been shown in the literature [44, 49], here, we provide the proof for completeness.
@@
The solution map $y^{*}_{\mu}:\mathcal{X}\rightarrow\mathcal{Y}$ is Lipschitz continuous. Suppose Assumptions 2.1, 2.2 hold and $\theta^{*}(x)$ is Lipschitz continuous map with constant $\mathbf{L}_{\theta}$. In particular, for any $x,\overline{x}\in\mathcal{X}$
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq\frac{L_{y\mu}}{\mu}\left\|x-\overline{x}\right\|,$
@@
$(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{\top}\big{[}\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\big{]}+\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq 0.$ (10)
@@
$(y-y^{*}_{\mu}(\overline{x}))^{\top}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))\leq 0,$ (11)
$(y-y^{*}_{\mu}(x))^{\top}\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\leq 0.$ (12)
@@
$(y^{*}_{\mu}(x)-y^{*}_{\mu}(\overline{x}))^{\top}\Big{(}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(x))\Big{)}\leq 0.$ (13)
@@
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{\top}\Big{(}\nabla_{y}\mathcal{L}_{\mu}(\overline{x},y^{*}_{\mu}(\overline{x}))-\nabla_{y}\mathcal{L}_{\mu}(x,y^{*}_{\mu}(\overline{x}))\Big{)}.$
@@
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq(y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x))^{\top}\Big{(}\nabla_{y}\Phi_{\mu}(\overline{x},\theta^{*}(\overline{x}),y^{*}_{\mu}(\overline{x}))-\nabla_{y}\Phi_{\mu}(x,\theta^{*}(x),y^{*}_{\mu}(\overline{x}))\Big{)}.$
@@
$\mu\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|^{2}\leq\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|(L_{yx}^{\Phi}\left\|x-\overline{x}\right\|+L_{y\theta}^{\Phi}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|),$
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq\frac{L_{yx}^{\Phi}}{\mu}\left\|x-\overline{x}\right\|+\frac{L_{y\theta}^{\Phi}}{\mu}\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|.$
@@
$\left\|y^{*}_{\mu}(\overline{x})-y^{*}_{\mu}(x)\right\|\leq(\frac{L_{yx}^{\Phi}+\mathbf{L}_{\theta}L_{y\theta}^{\Phi}}{\mu})\left\|x-\overline{x}\right\|.$
@@
Suppose Assumptions 2.1, 2.2 hold. The function $f_{\mu}(\cdot)$ is differentiable on an open set containing $\mathcal{X}$ and $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$ where $y^{*}_{\mu}(x)\triangleq\operatorname*{\arg\!\min}_{y\in\mathcal{Y}}\mathcal{L}_{\mu}(x,y)$. Moreover, $f_{\mu}$ has a Lipschitz continuous gradient with constant $\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu}$.
@@
From Danskins’s theorem [3] one can obtain $f_{\mu}(\cdot)$ is differentiable and $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$. Therefore, we have for any $x,x^{\prime}\in\mathcal{X}$,
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|=\left\|\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))-\nabla_{x}\mathcal{L}(x^{\prime},y^{*}_{\mu}(x^{\prime}))\right\|.$
@@
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-x^{\prime}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\left\|y^{*}_{\mu}(x)-y^{*}_{\mu}(x^{\prime})\right\|.$
@@
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-x^{\prime}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\frac{L_{y\mu}}{\mu}\left\|x-x^{\prime}\right\|,$
$\left\|\nabla f_{\mu}(x)-\nabla f_{\mu}(x^{\prime})\right\|\leq(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\left\|x-x^{\prime}\right\|.$
Next, in the subsequent lemmas, we establish the bound on the distance between the iterative solution $y_{k}$ and the optimal solution $y^{*}_{\mu}(x_{k})$. Moreover, by deriving a bound on the consecutive iterates of $y_{k}$, we demonstrate that the error between them can be controlled using parameter $\gamma$.
@@
Suppose Assumptions 2.1, 2.2 hold. Let $\{(x_{k},y_{k})\}_{k\geq 0}$ be the sequence generated by the Algorithm 1 or 2 with $\sigma_{k}=\sigma\leq\frac{2}{L_{yy}^{\Phi}+2\mu}$. Then for any $k\geq 0$,
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)},$
@@
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|\leq\frac{L^{\Phi}_{y\theta}}{\mu}\Big{(}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)}.$
@@
$\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma\|G_{k-1}^{y}-\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k-1})\|,$
@@
$\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}\|\theta_{k}-\theta^{*}(x_{k})\|.$
@@
$\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|$ $\leq\rho_{D}\|y_{k-1}-y^{*}_{\mu}(x_{k})\|+\sigma L^{\Phi}_{y\theta}(\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|x_{i}-x_{i+1}\right\|).$
@@
where $\Gamma_{1}\triangleq\frac{1}{1-\rho_{D}}\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|D_{\mathcal{X}}+\frac{\sigma L_{y\theta}^{\Phi}\beta D_{\mathcal{X}}}{(1-\rho_{D})(1-\beta)}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$ and $\Gamma_{2}\triangleq\frac{\gamma L_{y\mu}}{\mu(1-\rho_{D})}+\frac{\sigma\gamma\beta L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{(1-\beta)(1-\rho_{D})}=\mathcal{O}\left(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma\right)$. Furthermore, if $L^{\Phi}_{yy}=0$, in (15) we have that $\Gamma_{1}=\frac{\sigma L^{\Phi}_{y\theta}}{(1-\beta)}\|\theta_{0}-\theta^{*}(x_{0})\|D_{\mathcal{X}}=\mathcal{O}(\frac{\kappa_{g}L^{\Phi}_{y\theta}}{\mu})$ and $\Gamma_{2}=\frac{\gamma L_{y\mu}}{\mu}+\frac{\sigma\gamma\beta L^{\Phi}_{y\theta}\mathbf{L}_{\theta}}{(1-\beta)}=\mathcal{O}(\frac{\kappa_{g}^{2}L^{\Phi}_{y\theta}}{\mu}\gamma)$.
@@
$\quad+\gamma\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|\Big{)},$
where in the last inequality, we applied Lemma 4.3.
@@
$\quad+\rho^{k-i}_{D}\gamma\mathbf{L}_{\theta}\sum_{j=0}^{i-1}\beta^{i-j}\left\|s_{j}-x_{j}\right\|\Big{)},$
@@
Moreover, using Lemma 8.1 and the fact that $\|s_{k}-x_{k}\|\leq D_{\mathcal{X}}$, one can verify that for any $K\geq 1$,
@@
Finally, if $L^{\Phi}_{yy}=0$, then $\rho_{D}=0$, selecting $\sigma_{k}=\sigma=\frac{1}{\mu}$, we have
$\sum_{k=0}^{K-1}\left\|y_{k}-y^{*}_{\mu}(x_{k})\right\|\|s_{k}-x_{k}\|$ $\leq\frac{L^{\Phi}_{y\theta}}{\mu(1-\beta)}\|\theta_{0}-\theta^{*}(x_{0})\|D_{\mathcal{X}}+\frac{L^{\Phi}_{y\theta}\gamma\mathbf{L}_{\theta}\beta}{\mu(1-\beta)}\sum_{k=0}^{K-1}\|s_{k}-x_{k}\|^{2}.$
Note that $\mathbf{L}_{\theta}=\mathcal{O}(\kappa_{g})$, $L_{y\mu}=\mathcal{O}(\kappa_{g})$, $1-\rho_{D}=\mathcal{O}(\mu)$, and $1-\beta=\mathcal{O}(1/\kappa_{g})$.
@@
Suppose Assumptions 2.1, 2.2 hold. For any $k\geq 0$ we have
@@
where $\Gamma_{3}\triangleq\frac{2\Gamma_{2}}{1-\rho}+\frac{L_{y\mu}\gamma}{(1-\rho)\mu}$, and $\Gamma_{1},\Gamma_{2}$ are defined in Lemma 4.6 and $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$. Furthermore, if $L^{\Phi}_{yy}=0$, then $\rho_{D}=0$, and the bound attained in (4.7) can be rewritten as
$\left\|y_{k+1}-y_{k}\right\|$ $\leq\frac{2L_{y\theta}^{\Phi}}{\mu}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|\beta^{k}+\frac{2\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}}{\mu}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|+\frac{L_{y\mu}}{\mu}\gamma\left\|s_{k}-x_{k}\right\|.$
In this case, $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{3}L^{\Phi}_{y\theta}}{\mu}\gamma+\frac{\kappa_{g}^{2}L^{\Phi}_{y\theta}+\kappa_{g}L^{\Phi}_{yx}}{\mu}\gamma)$.
@@
Using the triangle inequality, one can observe that for any $k\geq 0$,
$\left\|y_{k+1}-y_{k}\right\|\leq\left\|y_{k+1}-y^{*}_{\mu}(x_{k+1})\right\|+\left\|y^{*}_{\mu}(x_{k})-y_{k}\right\|+\left\|y^{*}_{\mu}(x_{k+1})-y^{*}_{\mu}(x_{k})\right\|.$
Using Lemma 4.3 and 4.6, we can obtain
@@
Moreover, the second part of the lemma can be proved with a similar discussion as presented in the proof of Lemma 4.6, by employing the results in Lemma 8.1.